# VisuRiddles: Fine-grained Perception is a Primary Bottleneck for Multimodal Large Language Models in Abstract Visual Reasoning

### v1

@@ Abstract
Recent strides in multimodal large language models (MLLMs) have significantly advanced their performance in many reasoning tasks. However, Abstract Visual Reasoning (AVR) remains a critical challenge, primarily due to limitations in perceiving abstract graphics. To tackle this issue, we investigate the bottlenecks in current MLLMs and synthesize training data to improve their abstract visual perception. First, we propose VisuRiddles, a benchmark for AVR, featuring tasks meticulously constructed to assess models’ reasoning capacities across five core dimensions and two high-level reasoning categories. Second, we introduce the Perceptual Riddle Synthesizer (PRS), an automated framework for generating riddles with fine-grained perceptual descriptions. PRS not only generates valuable training data for abstract graphics but also provides fine-grained perceptual description, crucially allowing for supervision over intermediate reasoning stages and thereby improving both training efficacy and model interpretability. Our extensive experimental results on VisuRiddles empirically validate that fine-grained visual perception is the principal bottleneck and our synthesis framework markedly enhances the performance of contemporary MLLMs on these challenging tasks. Our code and dataset will be released at https://github.com/yh-hust/VisuRiddles
@@
Figure 1: Performance and analysis of MLLMs on AVR. (a) Most advanced MLLMs achieve limited accuracy on VisuRiddles, often close to random choice and far below human performance. (b) Model responses to abstract graphics and their perceptual descriptions show that once equipped with perceptual understanding, MLLMs can succeed in AVR task.
@@ 1 Introduction
Visual reasoning is a core component of human intelligence [1]. Enhancing this capacity in Multimodal Large Language Models (MLLMs) is key to bridging the gap to human-level performance [10, 11]. Recent advancements [4, 5, 6, 75, 9, 8, 7] in MLLMs have made significant progress in various general visual tasks, but still struggle with Abstract Visual Reasoning (AVR). As shown in Fig. 1(a), even advanced models such as Gemini2.5 Pro continue to face significant challenges in this task, showing a substantial gap compared to human performance. While such AVR tasks are relatively easy for humans, even the most advanced MLLMs struggle to solve them. It seems unexpected that models with strong performance in general understanding and reasoning tasks still fall short in handling such reasoning challenges. In fact, the core issue lies not in reasoning but in the fine-grained perception of complex visual structures. Unlike humans, who naturally process fine-grained visual structural cues, MLLMs often lack this perceptual ability needed for high-level abstract reasoning. As shown in Fig. 1(b), once the abstract graphics is reformulated into fine-grained perceptual descriptions, the model can produce accurate responses. This highlights that enhancing fine-grained perceptual capacity is crucial to unlocking MLLMs’ potential in AVR.
The significance of AVR tasks derives not only from their inherent complexity but also from their role in revealing the fundamental nature of human perception and cognition [62]. Therefore, abstract graphics serves as a prototypical and challenging paradigm for evaluating the reasoning capabilities of MLLMs [2, 3]. Cognitive research suggests that AVR requires the integration of multiple visual dimensions within abstract graphics, including numerosity, attributes, style, position, and spatial relations, which correspond respectively to sensitivity to quantity, object features, stylistic variation, relative placement, and spatial arrangement [63, 64, 65, 66]. Visual abstraction abilities across these five dimensions form the foundation of visual pattern understanding and serve as indispensable cognitive supports for high-level visual logical reasoning. A deficiency in perceiving the visual information of these dimensions can significantly limit a MLLM’s performance on AVR tasks.
Although solving abstract graphics problems involves multiple cognitive dimensions, benchmarks explicitly designed to evaluate them along these dimensions remain limited. To address this gap, we introduce VisuRiddles, a benchmark for evaluating AVR based on the five core dimensions and two high-level reasoning categories. VisuRiddles enables a more comprehensive assessment of abstract visual reasoning. However, most existing datasets offer only question–answer pairs, lacking the perceptual process supervision needed to bridge visual input and reasoning. To tackle this challenge, we further introduce an automated data synthesis framework that generates aligned samples with fine-grained perceptual descriptions, enabling granular supervision across the full perception-to-reasoning pipeline for more systematic learning and improved generalization.
@@
  • We introduce VisuRiddles, a comprehensive benchmark for AVR, enabling multi-dimensional evaluation of MLLMs across diverse reasoning dimensions.
  • We present the Perceptual Riddle Synthesizer, an automated framework designed to address the lack of intermediate supervision by producing structured AVR riddles enriched with fine-grained perceptual descriptions.
  • We conduct extensive experiments to demonstrate our key findings, including the reasoning limitations of current MLLMs, the effectiveness of our synthesized riddles, and the effective approach to addressing abstract visual problems.
@@ 2 Related Work
High-quality benchmarks are essential for driving progress in MLLMs. Earlier evaluation effort predominantly centered on basic visual comprehension tasks, leading to the development of extraction-oriented benchmarks [12, 13, 14, 15, 76, 76, 77, 16]. These benchmarks primarily assess a model’s ability to identify and align explicit visual elements. However, they depend heavily on surface-level cues, offering little assessment of an MLLM’s capacity for abstraction and logical reasoning. Therefore, researchers have expanded their focus on more holistic evaluation frameworks, proposing general-purpose benchmarks [17, 18, 19, 20, 78]. These benchmarks span multiple tasks, including visual understanding, mathematical reasoning, and knowledge-based QA, and are designed to assess the generalist capabilities of MLLMs. Although general-purpose benchmarks span a wide range of tasks, they tend to emphasize knowledge-based evaluation over visual logical reasoning. Moreover, due to their strong semantic dominance, MLLMs may directly rely on their inherent knowledge, making it difficult to disentangle and evaluate models’ true visual reasoning capabilities. To address these limitations, certain studies have shifted toward evaluating visual logical reasoning, leading to logic-oriented benchmarks such as PGM [24], RAVEN [21], I-RAVEN [22], RAVEN-FAIR [23], CVR [25], Raven IQ [26], ConceptARC [27], RPMs [28], MaRs-VQA [2] MathVista [29], GeoEval [30], MathVision[31], MathVerse [32], CMMath [33], OCR-Reasoning [78] and MV-MATH [34]. These benchmarks use structured diagrams, mathematical forms, and spatial layouts to assess analogical reasoning, compositionality, and pattern induction, signaling a move toward higher-level AVR. However, these benchmarks still exhibit notable limitations, particularly in their partial reliance on external knowledge and lack breadth in reasoning coverage.
2.2 Multimodal Large Language Models for Visual Reasoning.
@@
Advanced commercial LLMs, including the GPT [35], Gemini [36], and Claude [37] series, have exhibited impressive capabilities in multimodal understanding and generation. These advances have stimulated increasing interest and systematic investigation into MLLMs from both academia and industry. Early representative general-purpose MLLMs, such as LLaVA [38], Instruct-BLIP [39], Qwen-VL [40], Intern-VL [41], and mPLUG-Owl [42], perform well on general multimodal tasks but exhibit limitations in fine-grained visual perception. Subsequently, DeepSeek-VL [43], Monkey [8], InternVL 1.5 [44], TextMonkey [45], InternLM-XComposer2-4KHD [46], Qwen2-VL [47], and mPLUG-Owl2 [48] have introduced high-resolution visual perception mechanisms, establishing a more robust foundation for fine-grained visual understanding and reasoning. However, such improvements primarily expand the perceptual scope of MLLMs without substantially enhancing their reasoning capabilities. Therefore, inference-time scaling methods including LLaVA-CoT[49], BBA [50], R-CoT [51], and RedStar [52] have introduced and innovated upon inference-time scaling techniques such as chain-of-thought (CoT), significantly enhancing the visual reasoning capabilities of MLLMs. More recently, following the advances of Kimi K1.5 [53] and DeepSeek-R1 [54], a series of studies, such as R1-OneVision [3], LMM-R1 [55], MM-EUREKA [56], R1-V [57], Visual-RFT [58], VisualPRM [59], OThink-MR1 [60], and VLM-R1 [61], have employed reinforcement learning to develop adaptive reasoning policies and improve generalization in complex visual reasoning tasks.
@@
3 Method
In this section, we first describe the construction process of VisuRiddles, following a structured pipeline consisting of data collection, cleaning, and consolidation. We then summarize key benchmark statistics—including question and image counts, as well as answer distribution in Sec. 3.1. Subsequently, we introduce the Riddles Synthesis Framework, an automated approach for generating structurally diverse and perceptually annotated samples tailored for AVR in Sec. 3.2.
3.1 VisuRiddles Benchmark
Existing benchmarks provide limited coverage of the core reasoning dimensions required for AVR, To address this, VisuRiddles explicitly evaluates models across five key dimensions: a) Numerosity assesses the model’s ability to perceive and reason about quantity and distribution; b) Attribute evaluates the understanding of intrinsic visual features that determine structural semantics; c) Style tests the capacity to identify and generalize transformation-based visual patterns; d) Position reflects the ability to reason over the relative positions and layout of visual elements; e) Spatiality examines the understanding of three-dimensional structures, shape variations, and spatial transformations in abstract scenes. To extend the evaluation beyond low-level perceptual reasoning and address limitations in task diversity, structural complexity, and the inherent uncertainty introduced by single-choice formats, we additionally incorporate two high-level reasoning tasks that require analogical abstract reasoning and consistency-based logical reasoning, as represented by RAVEN and Sudoku, respectively. Finally, VisuRiddles also includes a subset, namely Other of AVR tasks involving planar shape composition and character-based semantic patterns. VisuRiddles provides a unified benchmark spanning basic to high-level reasoning, enabling comprehensive assessment of AVR, with representative samples shown in Fig. 2.
Collection. We manually collect a diverse set of visual reasoning problems from Chinese National Civil Service Examination, which form the five core categories of the benchmark. RAVEN and Sudoku samples are generated following the method1111https://github.com/WellyZhang/RAVEN1 and the protocols2222https://gitee.com/mxx11/sudoku2.
Following the structured pipeline, we initially collected 1,275 abstract visual reasoning problems, each accompanied by expert-level analysis and answer. After careful filtering and quality control, we construct VisuRiddles with 770ok samples in the basic categories and 200 samples in the high-level categories. Key statistics of VisuRiddles are summarized in Tab. 2. For the basic AVR tasks, each question is formulated as a single-choice problem, with answer option distributions of A (25.75%), B (24.75%), C (24.375%), and D (25.125%). In the high-level categories, RAVEN Reasoning and Sudoku Reasoning each account for 10% of the total data, and require models to generate exact symbolic outputs to be considered correct.
@@ 3.1 VisuRiddles Benchmark
Type Questions Number Question length Image Number Answer Formats
Sudoku Reasoning 100 48.3 100 Constraint-based grid
RAVEN Reasoning 100 70.7 200 Pattern match index
Numerical Reasoning 250 26.3 250 Single-choice
Stylistic Reasoning 117 26.3 117 Single-choice
Attribute Reasoning 97 27.0 97 Single-choice
Positional Reasoning 111 25.5 111 Single-choice
Spatial Reasoning 156 27.2 156 Single-choice
Other 69 30.3 69 Single-choice
Total 1000 32.5 1100
3.2 Perceptual Riddle Synthesizer
Recent MLLMs demonstrate strong general reasoning, but their limited performance on AVR is largely attributed to insufficient capacity of fine-grained perception. Most existing datasets lack intermediate perceptual description, providing only question-answer pairs. This prevents explicit modeling of the perception-to-reasoning process, leading to black-box inference, weak inductive capacity, and poor generalization. More critically, annotating such perceptual processes requires effort of expert-level staff members, further compounding this problem. To address this, we introduce Perceptual Riddle Synthesizer (PRS), a Riddles Synthesis Framework to provides abstract graphics with aligned perceptual description.
Figure 3 illustrates the pipeline of our PRS framework for AVR task. The pipeline consists of three main stages:
Rule Selection. The synthesis process begins with rule selection. We first extract representative reasoning patterns from real-world standardized tests, such as civil service visual logic exams. These high-level rules are categorized into positional, stylistic, numerical,attribute,spatial,raven and sudoku types, and further decomposed into finer-grained sub-rules, each representing a distinct visual transformation or logical operation (e.g., “simultaneous rotation and revolution”, or “black-white XOR operation”). This decomposition enables modular and interpretable synthesis, allowing each instance to target a specific reasoning skill.
Element Configuration. Once a sub-rule is selected, we proceed to define the visual components in a structured manner. This includes: (i) choosing a layout template (e.g., $3\times 33 × 3$ matrix, progression row, etc.), (ii) selecting or auto-generating icon elements with specific visual properties (e.g., shape, color, direction), and (iii) assigning a background grid. Elements are configured in a rule-consistent way: for example, a “positional” sub-rule may require translation across rows, while a “stylistic” sub-rule may apply logical operations to shape fill. Each configuration ensures the pattern remains visually interpretable and solvable.
Rendering and Annotating.
PRS produces a large-scale dataset spanning five core perceptual reasoning types and two high-level structured reasoning tasks. It supports 34 sub-rules across these categories and generates samples using 8 layout and 10k unique graphical icons. Each instance is created through rule-guided visual composition and rendered alongside structured perceptual annotations. The resulting dataset exhibits high diversity in both visual patterns and reasoning logic, as summarized in Tab. 2. It is worth noting that the synthetic tasks are designed to strengthen perceptual abstraction rather than replicate real-world reasoning difficulty. Consequently, their complexity is deliberately kept tractable to focus on providing structured perceptual supervision.
@@ 3.2 Perceptual Riddle Synthesizer
Figure 3: Overview of the PRS framework. (a) A unified pipeline for generating abstract graphics with fine-grained perceptual descriptions across different rule categories. (b) Visualization of synthesized riddles based on positional rule and stylistic rule.
Table 2: Configuration details of PRS across seven reasoning categories.
Rule Sub-rule Layout Template Icon Sample
Numerical 5 2 1 5k 10K
Stylistic 6 3 5 5k 10K
Attribute 2 2 1 5k 10K
Positional 8 4 7 5k 10K
Spatial 5 2 13 5k 10K
Sudoku 1 1 4 9 10K
RAVEN 7 5 5 10k 20K

### v3

@@ Abstract
Recent strides in multimodal large language models (MLLMs) have demonstrated significant progress in many reasoning tasks, but they still fail in Abstract Visual Reasoning (AVR) tasks. Our experimental findings indicate that the core bottleneck lies not only in the reasoning capabilities of MLLMs but more critically in their absence of fine-grained perception. To address this issue, we present VisuRiddles, a dedicated resource for AVR research. It consists of (i) a benchmark, collected from real-world data, for the systematic evaluation of MLLMs’ AVR capabilities, and (ii) a synthesizer, which automatically generates AVR instances enriched with perceptual descriptions and reasoning chains, enabling supervised training and deeper investigation. Building on VisuRiddles, we propose a two-stage training paradigm that progressively enhances perceptual ability and strengthens reasoning, producing the Perception-Augmented Visual Reasoner (PAVR). Experiments demonstrate that PAVR unifies perception and reasoning, substantially outperforming both open-source and commercial MLLMs, thereby underscoring fine-grained perception as the primary bottleneck in AVR. Our code and dataset will be released at https://github.com/yh-hust/VisuRiddles
@@
Figure 1: Performance and analysis of MLLMs on AVR. (a) Most advanced MLLMs achieve limited accuracy on VisuRiddles, often close to random choice and far below human performance. (b) Model responses to abstract graphics and their perceptual descriptions show that once equipped with perceptual capability, MLLMs can succeed in AVR task.
@@ 1 Introduction
Reasoning is a core component of human intelligence [1]. Enhancing this capability in Multimodal Large Language Models (MLLMs) is key to bridging the gap to human-level performance [10, 11]. Recent advancements [4, 5, 6, 75, 9, 8] in MLLMs have made significant progress in various general visual tasks, but still struggle with Abstract Visual Reasoning (AVR) tasks [84]. As shown in Fig. 1(a), even advanced models such as Gemini2.5 Pro continue to face significant challenges in this task, showing a substantial gap compared to human performance. While such AVR tasks are relatively easy for humans, even the most advanced MLLMs struggle to solve them. It unexpected that models with strong performance in general understanding and reasoning tasks still fall short in handling such reasoning challenges. In fact, the challenges of AVR primarily stem from two aspects: fine-grained perception and reasoning. Compared to the recent advances in enhancing reasoning capabilities, the ability to perceive subtle visual structures, such as position, style, and attribute, has been largely overlooked in existing studies [80, 79]. Unlike humans, which naturally process such fine-grained visual cues, MLLMs often lack the perceptual ability required for abstract graphics understanding [85, 2]. As illustrated in Fig. 1(b), once abstract graphics are reformulated into structured perceptual descriptions, the model is able to generate accurate responses. This highlights that strengthening perceptual capability is a crucial factor for advancing AVR. Therefore, genuine progress in enabling MLLMs to master AVR requires addressing both perception and reasoning, with special emphasis on enhancing the often-overlooked perceptual capability.
To advance the study of AVR tasks, we introduce the VisuRiddles benchmark, which is primarily derived from real riddles and enables objective evaluation of MLLMs’ performance. Since existing datasets lack fine-grained perceptual annotations, we further design the VisuRiddles Synthesizer, which automatically generates AVR instances enriched with structured perceptual descriptions. Leveraging this synthesized data, we perform Supervised Fine-Tuning (SFT) to enhance the model’s ability to perceive fine-grained visual cues, thereby equipping it with the perceptual foundation necessary for AVR. Although SFT substantially improves the perceptual capability of MLLMs, they still face persistent bottlenecks in AVR, including errors in perceptual strategy (e.g., difficulty in determining whether the appropriate interpretive cue lies in the figure’s symmetry axis or in its right-angle structures) and limited reasoning ability on more difficult instances. Therefore, we introduce Reinforcement Learning (RL), which leverages the enhanced perceptual capability to guide models toward more reliable perceptual grounding and further strengthen their reasoning ability.
Building upon these components, we develop the Perception-Augmented Visual Reasoner (PAVR), which unifies enhanced perceptual capability with improved reasoning ability to address the dual challenges of AVR. Experimental results show that PAVR significantly outperforms advanced commercial models on AVR tasks, highlighting a systematic solution to AVR tasks.
@@
(1) We introduce VisuRiddles, which comprises a benchmark derived from real riddles for objectively evaluating MLLMs on AVR tasks, and a synthesizer that automatically generates AVR instances enriched with structured perceptual descriptions.
(2) We leverage synthesized data for SFT to enhance fine-grained perceptual capability for AVR tasks, and further adopt RL to improve reasoning ability and stabilize perceptual strategy selection, resulting in the Perception-Augmented Visual Reasoner.
(3) We conduct extensive experiments to demonstrate our key findings, including the reasoning limitations of current MLLMs, the effectiveness of our synthesized riddles, and the effective approach to addressing abstract visual problems.
@@ 2.1 Multimodal Benchmarks
High-quality benchmarks are essential for driving progress in MLLMs. Earlier evaluation effort predominantly centered on basic visual comprehension tasks, leading to the development of extraction-oriented benchmarks [12, 13, 14]. These benchmarks primarily assess a model’s ability to identify and align explicit visual elements. However, they depend heavily on surface-level cues, offering little assessment of an MLLM’s capability for abstraction and logical reasoning. Therefore, general-purpose benchmarks [17, 19, 86, 20, 18] are widely focused on. These benchmarks span multiple tasks, such as visual understanding and mathematical reasoning, and are designed to assess the generalist capabilities of MLLMs. However, they tend to emphasize knowledge-based evaluation over visual logical reasoning. Moreover, due to their strong semantic dominance, MLLMs may directly rely on their inherent knowledge, making it difficult to disentangle and evaluate models’ true visual reasoning capabilities. Therefore, some studies have shifted toward evaluating visual logical reasoning, leading to logic-oriented benchmarks such as RAVEN [21], I-RAVEN [22], RAVEN-FAIR [23], CVR [25], RPMs [28], MaRs-VQA [2], MathVerse [32], MARVEL [79], PuzzleVQA [83], VisCogBench [2], VisuLogic [80], VisualPuzzle [81], VisualSphinx [82] and MV-MATH [34]. These benchmarks use structured diagrams, mathematical forms, and spatial layouts to assess analogical reasoning, compositionality, and pattern induction, signaling a move toward higher-level AVR. However, these benchmarks still exhibit notable limitations, particularly in their partial reliance on external knowledge and lack breadth in reasoning coverage.
2.2 Multimodal Large Language Models for Visual Reasoning
@@
Advanced commercial LLMs, including the GPT [35], Gemini [36], and Claude [37] series, have exhibited impressive capabilities in multimodal understanding and generation. These advances have spurred systematic research on MLLMs in both academia and industry. Early representative general-purpose MLLMs, such as LLaVA [38], Instruct-BLIP [39], Qwen-VL [40], and Intern-VL [41], perform well on general multimodal tasks but exhibit limitations in fine-grained visual perception. Subsequently, DeepSeek-VL [43], Monkey [8], TextMonkey [45], InternLM-XComposer2-4KHD [46], Qwen2-VL [47] have introduced high-resolution visual perception mechanisms, establishing a more robust foundation for fine-grained visual understanding and reasoning. However, such improvements primarily expand the perceptual scope of MLLMs without substantially enhancing their reasoning capabilities. Therefore, inference-time scaling methods including MMVP [85], LLaVA-CoT [49], BBA [50], R-CoT [51], and RedStar [52] have introduced and advanced inference-time scaling techniques such as chain-of-thought (CoT), significantly enhancing the visual reasoning capabilities of MLLMs. More recently, following the advances of Kimi K1.5 [53] and DeepSeek-R1 [54], a series of studies, such as R1-OneVision [3], LMM-R1 [55], MM-EUREKA [56], R1-V [57], Visual-RFT [58], VisualPRM [59], OThink-MR1 [60], and VLM-R1 [61], have employed reinforcement learning to develop adaptive reasoning policies and improve generalization in complex visual reasoning tasks.
3 VisuRiddles
In this section, we introduce VisuRiddles, a innovative and comprehensive resource for AVR research. It consists of the VisuRiddles Benchmark (Sec. 3.1), built from real-world problems to objectively evaluate MLLMs’ AVR performance, and the VisuRiddles Synthesizer (Sec. 3.2), designed to generate diverse perceptually annotated AVR instances for model training.
3.1 VisuRiddles Benchmark
Unlike RAVEN [21] and PuzzleVQA [83], which focus on general visual reasoning tasks, we target the most challenging part of AVR tasks. Thus, the VisuRiddles benchmark has been constructed to evaluate MLLMs. It explicitly evaluates models across five key dimensions: a) Numerosity assesses the model’s ability to perceive and reason about quantity and distribution; b) Attribute evaluates the understanding of intrinsic visual features that determine structural semantics; c) Style tests the capability to identify and generalize transformation-based visual patterns; d) Position reflects the ability to reason over the relative positions and layout of visual elements; e) Spatiality examines the understanding of three-dimensional structures, shape variations, and spatial transformations in abstract graphics. To extend the evaluation beyond low-level perceptual reasoning and address limitations in task diversity, structural complexity, and the inherent uncertainty introduced by single-choice formats, we additionally incorporate two high-level reasoning tasks that require analogical abstract reasoning and consistency-based logical reasoning, as represented by RAVEN and Sudoku, respectively. Finally, VisuRiddles also includes a subset, namely Other of AVR tasks involving planar shape composition and character-based semantic patterns. VisuRiddles provides a unified benchmark spanning basic to high-level reasoning, enabling comprehensive assessment of AVR, with representative samples shown in Fig. 2. We construct the VisuRiddles benchmark through a structured pipeline comprising Collection, Cleaning, and Consolidation. We provide the construction details in Appendix A.
@@ 3.1 VisuRiddles Benchmark
We initially collected 1,275 AVR problems, each accompanied by expert-level analysis and answer. After careful filtering and quality control, we construct VisuRiddles with 800 samples in the basic categories and 200 samples in the high-level categories. Key statistics of VisuRiddles are summarized in Tab. 1. For the basic AVR tasks, each question is formulated as a single-choice problem, with answer option distributions of A (25.75%), B (24.75%), C (24.375%), and D (25.125%). In the high-level categories, RAVEN Reasoning and Sudoku Reasoning each account for 10% of the total data, and require models to generate exact symbolic outputs to be considered correct.
@@
Type Questions Number Question length Answer Formats
Sudoku 100 48.3 Constraint-based grid
RAVEN 100 70.7 Single-choice
Numerical 250 26.3 Single-choice
Stylistic 117 26.3 Single-choice
Attribute 97 27.0 Single-choice
Positional 111 25.5 Single-choice
Spatial 156 27.2 Single-choice
Other 69 30.3 Single-choice
Total 1000 32.5 -
@@
Figure 3: Overview of the VisuRiddles Synthesizer. (a) A unified pipeline for generating abstract graphics with fine-grained perceptual descriptions. (b) Visualization of synthesized riddles based on positional rule and stylistic rule.
3.2 VisuRiddles Synthesizer
Recent MLLMs demonstrate strong general reasoning, but their limited performance on AVR is largely attributed to insufficient capability of fine-grained perception for abstract graphics. Most existing datasets lack intermediate perceptual description, providing only question-answer pairs [80, 79, 81]. This prevents explicit modeling of the perception-to-reasoning process, leading to black-box inference, weak inductive capability, and poor generalization. More critically, existing commercial models struggle to annotate perceptual processes effectively, and manual labeling requires significant expert effort, both of which further exacerbate the issue. To address this, we introduce VisuRiddles Synthesizer, a Riddles Synthesis Framework to provides abstract graphics with aligned fine-grained perceptual description.
Figure 3 illustrates the pipeline of our VisuRiddles Synthesizer for AVR task. The VisuRiddles Synthesizer pipeline comprises two stages: Riddles Construction and API Labeling. The former generates abstract visual instances with aligned perceptual descriptions, while the latter generates reasoning chains based on these descriptions. Details and representative examples are provided in Appendix B.2 and Appendix B.3.
The VisuRiddles Synthesizer generates data that spans five core perceptual reasoning types and two high-level reasoning tasks. The detailed configurations of VisuRiddles Synthesizer are provided in Table 5. The VisuRiddles Synthesizer not only provides riddles’ visual instances but also structured perceptual descriptions, thereby enhancing the model’s fine-grained perceptual ability for AVR tasks. It is noteworthy that the riddles instances generated by the VisuRiddles Synthesizer focus on improving fine-grained perception, rather than reasoning. Therefore, their reasoning difficulty is deliberately kept lower than that of real-world riddles.
4 Perception-Augmented Visual Reasoner
To tackle the challenges of AVR, we introduce the Perception-Augmented Visual Reasoner (PAVR). As illustrated in Fig. 4, PAVR is developed through a two-stage training paradigm. In the first stage, SFT equips the model with the ability to capture fine-grained visual cues in abstract graphics. Building upon this, the second stage employs RL to effectively integrate these visual cues into stable reasoning strategies for solving complex tasks, resulting in PAVR. These two stages are complementary: SFT enables the model to see clearly, while RL guides it to reason reliably.
Figure 4: Overview of PAVR. (i) Baseline: Incorrect perception leads to incorrect results. (ii) PAVR-SFT, which is trained on synthesized data to enhance fine-grained perceptual ability, can accurately understand visual content in riddles but fails in pattern identification. (iii) PAVR, which builds upon PAVR-SFT with reinforcement learning, can effectively recognize patterns and derive the correct answer.
Perceptual Augment via Supervised Fine-Tuning. The lack of fine-grained perceptual ability remains a major bottleneck for MLLMs on AVR tasks, yet existing real-world datasets typically provide only single-choice answer and rarely provide explicit perceptual annotations to supply sufficient supervision. To address this limitation, we employ the VisuRiddles Synthesizer described in Sec. 3.2 to generate AVR instances across seven categories, each enriched with structured perceptual descriptions. These synthesized examples serve as training data for SFT, enhancing the model’s perceptual ability and enabling it to capture fine-grained cues in abstract graphics. This establishes a reliable foundation for subsequent reasoning optimization through reinforcement learning.
Reasoning Optimization via Reinforcement Learning. Although SFT on synthesized data improves fine-grained perceptual ability, it still faces limitations in AVR, including unstable perceptual strategy selection (e.g., axis vs. angles) and insufficient reasoning on complex tasks. Therefore, we apply Group Relative Policy Optimization (GRPO) [54] to endow the model with more reliable perceptual grounding and stronger reasoning ability. The reward design includes: (i) Answer Reward, giving 1 for correct answers and 0 otherwise, and (ii) Format Reward, incentivizing outputs that match the required template: <think> … </think> <answer> … </answer>. The training data for GRPO is likewise generated by the VisuRiddles Synthesizer. This process enables the model to achieve more reliable perceptual grounding and stronger reasoning capabilities.